OpenStudy (anonymous):

In 1995, the life expectancy of males in a certain country was 62.1 years. In 1999, it was 65.7 years. Let E represent the life expectancy in year t and let t represent the number of years since 1995. what is the linear function? Use the function to predict the life expectancy of males in 2006?

OpenStudy (anonymous):

Use a simple rise over run. Where rise is equal to the # of years in 1999 minus the years in 1995 (65.7-62.1) and run is equal to 1999 minus 1995 = 4. thus, slope is equal to 3.6/4, and set the y intercept equal to 62.1 (at what we will call time t=0). Thus we get the equation E(t) = .9x+62.1, plug in 11 into x to get the value at 2006

OpenStudy (radar):

E(t)=.9t+62.1 t in years

OpenStudy (radar):

I was hesitant in answering as being 72 I had to be sure I was alive before proceeding. Tsizzle gave a more detailed and a good answer.

OpenStudy (radar):

Did you plug in the 11 for x and get an answer for 2006?

OpenStudy (anonymous):

no not sure how to do that.. very puzzled with this problem.. can u help????

OpenStudy (radar):

O.K. Use the Tsizzle formula: E=(.9)(11)+62.1 E=9.9+62.1 E=72 Great 72, I may live a much longer life now. lol 72 yrs of life expectancy in 2006

OpenStudy (anonymous):

lol..thanks for the help.

OpenStudy (anonymous):

sorry, I didn't realize I used both x and t in my formula, sorry for the confusion